### v1

@@ Abstract
We introduce kernel density machines (KDM), a novel density ratio estimator in a reproducing kernel Hilbert space setting. KDM applies to general probability measures on countably generated measurable spaces without restrictive assumptions on continuity, or the existence of a Lebesgue density. For computational efficiency, we incorporate a low-rank approximation with precisely controlled error that grants scalability to large-sample settings. We provide rigorous theoretical guarantees, including asymptotic consistency, a functional central limit theorem, and finite-sample error bounds, establishing a strong foundation for practical use. Empirical results based on simulated and real data demonstrate the efficacy and precision of KDM.
Keywords: density ratio estimation, reproducing kernel Hilbert space (RKHS), low-rank approximation, finite-sample guarantees
@@ 1 Introduction
Estimating the density ratio $g_{\star}\coloneqq\frac{d\mathbb{Q}}{d\mathbb{P}}$ of the probability measure $\mathbb{Q}$ with respect to $\mathbb{P}$ is central to many tasks in machine learning, statistics, and applied sciences, such as domain adaption and transfer learning (Pan and Yang, 2010), anomaly detection (Cover and Hart, 1967), causal inference (Imbens and Rubin, 2015), conditional distribution estimation (Racine, 2008; Muandet et al., 2017), and generative modeling. Extant approaches to density ratio estimation often assume that both $\mathbb{P}$ and $\mathbb{Q}$ are continuous with respect to the Lebesgue measure, restricting their applicability to fully continuous data. In contrast, we introduce a kernel-based method that applies to general probability measures on countably generated measurable spaces, encompassing settings with discrete, continuous, mixed, or even singular distributions: kernel density machines (KDM). The parsimony of KDM with respect to the underlying assumptions significantly expands the applicability of density ratio estimation.
@@
We model $g_{\star}$ as the sum of an exogenous prior plus a function in a reproducing kernel Hilbert space (RKHS), which, while possessing attractive theoretical properties, is known to pose computational problems with large data sets. To ensure scalability in the sample size and also higher-dimensional applications, we therefore introduce a low-rank approximation that retains critical information while reducing computational costs. The approximation error associated with the low-rank density ratio is sharply controlled by the error of a low-rank approximation of the kernel matrix pertaining to the RKHS and the data points. This combination of flexibility and computational efficiency makes our method applicable to various real-world scenarios, from structured data in financial modeling, to complex conditional distributions in scientific research.
@@
We provide rigorous theoretical results, including asymptotic consistency, a functional central limit theorem, and finite-sample bounds for our estimator, and in particular also its low-rank approximation. These guarantees make KDM reliable across different sample sizes and data configurations. As an application, we obtain statistical tests of the prior density ratio and recover the kernel two-sample test of Gretton et al. (2012) as a special case. As a second application, we obtain a novel, consistent estimator of conditional distributions. We illustrate the applicability of KDM with both simulated and real data, including tests for independence and conditional distributions. [^4] An elementary reference implementation can be downloaded from KDM.R. The empirical results suggest KDM to be useful, in particular for larger data sets and higher dimensions than usually considered in RKHS and local nonparametric kernel density estimators (Li and Racine, 2006).
@@
Our approach builds on recent advances in kernel-based density ratio estimation, but extends beyond existing methods by applying to general probability measures. Sugiyama et al. (2012, 2008) and Filipović et al. (2024) are closest to our approach, but are restricted to Lebesgue densities and do not provide theoretical guarantees, respectively. Another related work is Zellinger et al. (2023), who provide optimal convergence rates of density ratio estimators under rather strong regularity assumptions (source conditions), as used in inverse problems and learning theory. This is in contrast to our more agnostic framework. They also do not discuss computational scalability, as we do with low-rank approximations. Another strand of literature, including (Song et al., 2008; Klebanov et al., 2021; Park and Muandet, 2020), estimates conditional expectations directly via conditional mean embeddings (CME). This is a powerful and empirically widely used technique. But there is no obvious way how to recover density ratio estimates from CMEs.
@@
The remainder of this paper is organized as follows. Section 2 formalizes the density ratio estimation problem for general probability measures and introduces our RKHS-based estimator. Section 3 presents theoretical guarantees, including consistency and finite-sample bounds. Section 4 discusses the low-rank approximation approach and its computational advantages. In Section 5, we apply our method to empirical tasks in domain adaptation, anomaly detection, causal inference, and high-dimensional conditional distribution estimation. Section 6 illustrates the applicablity of KDM in hypothesis testing and conditional distribution estimation for both simulated and real data. Section 7 concludes and points out future directions. The appendix contains additional results and all proofs.
2 Density Ratio Learning Problem
@@
We denote the density ratio (Radon–Nikodym derivative) by $g_{\star}\coloneqq\frac{d{\mathbb{Q}}}{d{\mathbb{P}}}$.
The goal of this paper is to learn the true density ratio $g_{\star}$ under minimal assumptions from samples of ${\mathbb{P}}$ and ${\mathbb{Q}}$. To this end, we assume that
@@
see, e.g., Steinwart and Scovel (2012, Section 2). Given (2), we can represent the hypothesis density ratio as
@@
The additive splitting (4) is motivated by the fact that $g_{\star}$ itself might not be an element of ${\mathcal{H}}$ while $g_{\star}-p_{\star}$ is. To see the relevance of this approach, consider the example where ${\mathcal{Z}}={\mathbb{R}}$, $g_{\star}=1$ and $k$ is a Gaussian kernel. Indeed, this Gaussian RKHS does not contain the constant functions (in fact, it contains no polynomials, see Minh (2010)). In this example, one could thus choose the constant prior $p_{\star}=1$. In general, we neither assume that $p_{\star}$ is positive nor that it integrates to 1 with respect to ${{\mathbb{P}}}$. Hence, $p_{\star}=0$ is a possible choice. Notably, our specification of the hypothesis density ratio (4) is novel to the literature, as it is more flexible and requires only the milder assumption (2), in contrast to the stronger assumptions typically employed in previous work. For a more detailed comparison, see Section 5.2.3 below.
@@
The error function ${\mathcal{E}}(h)$, and the expression in (8), contain the unobservable population density ratio $g_{\star}$ and therefore are not readily available for estimations based on samples of ${\mathbb{P}}$ and ${\mathbb{Q}}$. We solve this problem by using the identity $J_{\mathbb{P}}^{\ast}g_{\star}=J_{\mathbb{Q}}^{\ast}1$ justified by the following lemma.
@@
To assess how close $p_{\star}+J_{\mathbb{P}}h_{\lambda}$ is to the true density ratio $g_{\star}$, we consider the orthogonal decomposition of the squared population error
@@ 4 Low-rank Approximation
The following classic result asserts that Algorithm 1 works, see, e.g., Horn and Zhang (2005) or Harbrecht et al. (2012). A novelty of Algorithm 1 lies in the computation of the biorthogonal matrix $\bm{R}$ at no additional complexity along with $\bm{L}$. For completeness, we provide a self-contained proof in Appendix A.
@@ 5.1 Hypothesis Testing
As a first application, we obtain statistical tests of the hypothesis that the true equals the prior density ratio,
@@
For $p_{\star}=1$, the hypothesis (38) is equivalent to ${\mathbb{Q}}={\mathbb{P}}$, which is the same hypothesis as studied in Pfister et al. (2017). Then the sample variable $u_{\lambda}$ in (82) is the scaled difference $S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{1}$ between the sample kernel mean embeddings of ${\mathbb{Q}}$ and ${\mathbb{P}}$. As a corollary of Theorem 5.1 we thus recover the kernel two-sample test of Gretton et al. (2012) as a special case, connecting their maximum mean discrepancy approach (Gretton et al., 2012, Theorem 5) with the Hilbert–Schmidt independence criterion (Pfister et al., 2017, Proposition 1).
@@ 5.2.1 Derivation of the Estimator
For any hypothesis density ratio (4) with $h\in{\mathcal{H}}$, we thus obtain the corresponding model of the true conditional expectation ${\mathbb{E}}[f(Y)\mid X=x]=\int_{\mathcal{Y}}f(y){\mathbb{P}}_{Y\mid X=x}(dy)$ given by
@@
for any $f\in L^{2}_{{\mathbb{P}}_{Y}}$. Here we estimate the ${\mathbb{P}}_{Y}$-integral in (49) by the empirical counterpart given by (50), for an auxiliary i.i.d. sample $\bar{y}_{1},\dots,\bar{y}_{\bar{n}}$ of ${\mathbb{P}}_{Y}$. The last approximation in (51) is practically motivated such that we integrate $f$ with respect to a bona-fide conditional distribution. [^12] The density ratio $g_{\star}$ satisfies the implicit structural properties $g_{\star}(x,y)\geq 0$ ${\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$-a.s., $\int_{\mathcal{Y}}g_{\star}(x,y){\mathbb{P}}_{Y}(dy)=1$ ${\mathbb{P}}_{X}$-a.s., and $\int_{\mathcal{X}}g_{\star}(x,y){\mathbb{P}}_{X}(dx)=1$ ${\mathbb{P}}_{Y}$-a.s.
The estimator $\widetilde{h}_{\lambda}$ in (49) requires i.i.d. samples $z_{{\mathbb{P}},1},\dots,z_{{\mathbb{P}},n}$ of ${\mathbb{P}}\coloneqq{\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$ and $z_{{\mathbb{Q}},1},\dots,z_{{\mathbb{Q}},n}$ of ${\mathbb{Q}}\coloneqq{\mathbb{P}}_{(X,Y)}$. If we could only sample from the joint distribution, we would consider an i.i.d. sample $(x_{1},y_{1}),\dots,(x_{3n},y_{3n})$ of ${\mathbb{P}}_{(X,Y)}$ of size $3n$ and set
@@ 5.2.3 Comparison with Related Literature
In view of the decomposition (48) it is natural to compare conditional expectations (46) calculated from a hypothesis density ratio (4) to the conditional mean embedding introduced in Song et al. (2009). To facilitate this comparison, we assume in this subsection that ${\mathcal{H}}={\mathcal{H}}_{\mathcal{X}}\otimes{\mathcal{H}}_{\mathcal{Y}}$ is the tensor product of two separable RKHS ${\mathcal{H}}_{\mathcal{X}}$ and ${\mathcal{H}}_{\mathcal{Y}}$ with measurable kernels $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$ on ${\mathcal{X}}$ and ${\mathcal{Y}}$, respectively. In line with (3), we also assume that $\sup_{x\in{\mathcal{X}}}k_{\mathcal{X}}(x,x)<\infty$ and $\sup_{y\in{\mathcal{Y}}}k_{\mathcal{Y}}(y,y)<\infty$, so that the canonical embeddings $J_{{\mathbb{P}}_{X}}\colon{\mathcal{H}}_{\mathcal{X}}\to L^{2}_{{\mathbb{P}}_{X}}$, $J_{{\mathbb{P}}_{Y}}\colon{\mathcal{H}}_{\mathcal{Y}}\to L^{2}_{{\mathbb{P}}_{Y}}$ are Hilbert–Schmidt operators. We then have $k\big{(}(x_{1},y_{1}),(x_{2},y_{2})\big{)}=k_{\mathcal{X}}(x_{1},x_{2})k_{\mathcal{Y}}(y_{1},y_{2})$, and in view of (47) we can identify the operators $J_{{\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}}=J_{{\mathbb{P}}_{X}}\otimes J_{{\mathbb{P}}_{Y}}$.
@@
Our approach, via the density ratio $g_{\star}$, overcomes the above limitations of the traditional conditional mean embedding approach. First, our construction of $\mu_{Y\mid X=x}$ is more elementary and feasible under verifiable technical assumptions, such as (1) and (2). Second, our hypothesis model (4) of $g_{\star}$ yields the candidate conditional distribution ${\mathbb{P}}_{Y\mid X=x}$ directly using identity (45). Third, we thus compute conditional expectations (46) directly for all test functions $f\in L^{2}_{{\mathbb{P}}_{Y}}$, and not only for $f\in{\mathcal{H}}_{\mathcal{Y}}$.
@@
Also the setup in Suzuki et al. (2009); Sugiyama (2013) is restricted to continuous variables with a Lebesgue density. In fact, they assume the assumptions of Lemma 5.3 hold for the Lebesgue measures $\mu_{\mathcal{X}}(dx)=dx$ and $\mu_{\mathcal{Y}}(dy)=dy$. In this case, the “density ratio $w(x,y)$” in Suzuki et al. (2009) is equal to our density ratio $g_{\star}(x,y)$. Their estimator of the density ratio is based on a sampling scheme, similar to Filipović et al. (2024), which is different from ours. As a consequence, they do not provide finite sample guarantees nor a central limit theorem.
There is a vast literature on (semi-)parametric distributional regression approaches, which includes structured additive models, see Rügamer et al. (2023) for a recent overview. In contrast, our approach is nonparametric functional analytic and learns the density ratio in a reproducing kernel Hilbert space, which comes with asymptotic theory and finite sample guarantees. Isotonic distribution regression in Henzi et al. (2021) is to date specialized on univariate response variables and conditional distributions, conformal predictive distributions from Vovk et al. (2019) are based on regression problems. Shen and Meinshausen (2024) estimates conditional distributions only.

### v2

@@ Abstract
We introduce kernel density machines (KDM), a nonparametric estimator of a Radon–Nikodym derivative, based on reproducing kernel Hilbert spaces. KDM applies to general probability measures on countably generated measurable spaces under minimal assumptions. For computational efficiency, we incorporate a low-rank approximation with precisely controlled error that grants scalability to large-sample settings. We provide rigorous theoretical guarantees, including asymptotic consistency, a functional central limit theorem, and finite-sample error bounds, establishing a strong foundation for practical use. Empirical results based on simulated and real data demonstrate the efficacy and precision of KDM.
Keywords: Radon–Nikodym derivative estimation, reproducing kernel Hilbert space (RKHS), low-rank approximation, finite-sample guarantees
@@ 1 Introduction
Estimating the Radon–Nikodym derivative $g_{\star}\coloneqq\frac{d\mathbb{Q}}{d\mathbb{P}}$ of the probability measure $\mathbb{Q}$ with respect to $\mathbb{P}$ is central to many tasks in machine learning, statistics, and applied sciences, such as domain adaption and transfer learning (Pan and Yang, 2010), anomaly detection (Cover and Hart, 1967), causal inference (Imbens and Rubin, 2015), conditional distribution estimation (Racine, 2008; Muandet et al., 2017), and generative modeling. We refer to $g_{\star}$ as the density of ${\mathbb{Q}}$ relative to ${\mathbb{P}}$ throughout, following standard usage in probability theory. [^4] This terminology follows the convention of (Jacod and Shiryaev, 1987, Chapter III.3). Extant approaches often assume that both $\mathbb{P}$ and $\mathbb{Q}$ admit Lebesgue densities, in which case $g_{\star}$ is commonly referred to as the density ratio, thereby limiting their applicability to purely continuous distributions. In contrast, we introduce a kernel-based method that applies to general probability measures on countably generated measurable spaces, encompassing settings with discrete, continuous, mixed, or even singular distributions: kernel density machines (KDM). The parsimony of KDM with respect to the underlying assumptions significantly expands its versatility with respect to applications.
@@
We model $g_{\star}$ as the sum of an exogenous prior plus a function in a reproducing kernel Hilbert space (RKHS), which, while possessing attractive theoretical properties, is known to pose computational problems with large data sets. To ensure scalability in the sample size and also higher-dimensional applications, we therefore introduce a low-rank approximation that retains critical information while reducing computational costs. The approximation error associated with the low-rank estimator is sharply controlled by the error of a low-rank approximation of the kernel matrix pertaining to the RKHS and the data points. This combination of flexibility and computational efficiency makes our method applicable to various real-world scenarios, from structured data in financial modeling, to complex conditional distributions in scientific research.
@@
We provide rigorous theoretical results, including asymptotic consistency, a functional central limit theorem, and finite-sample bounds for our estimator, and in particular also its low-rank approximation. These guarantees make KDM reliable across different sample sizes and data configurations. As an application, we recover the kernel two-sample test of Gretton et al. (2012) as a special case of our framework. As a second application, we obtain a novel, consistent estimator of conditional distributions. We illustrate the applicability of KDM with both simulated and real data, including tests for independence and conditional distributions. [^5] An elementary reference implementation can be downloaded from KDM.R. The empirical results suggest KDM to be useful, in particular for larger data sets and higher dimensions than usually considered in RKHS and local nonparametric kernel density estimators (Li and Racine, 2006).
@@
Our approach builds on recent advances in kernel-based density estimation, but extends beyond existing methods by applying to general probability measures. Sugiyama et al. (2012, 2008) and Filipovic et al. (2025) are related to our approach, but are either restricted to ratios of Lebesgue densities or do not provide the same theoretical guarantees and low-rank approximation error estimates as we do, respectively. Another related work is Nguyen et al. (2024); Zellinger et al. (2023), who provide optimal convergence rates under rather strong regularity assumptions (source conditions), as used in inverse problems and learning theory. This is in contrast to our more agnostic framework. They also do not discuss computational scalability, as we do with low-rank approximations. Another strand of literature, including (Song et al., 2008; Klebanov et al., 2021; Park and Muandet, 2020), estimates conditional expectations directly via the conditional mean embedding (CME). This is a powerful and empirically widely used technique, but there is no obvious way how to recover the density from the CME.
@@
The remainder of this paper is organized as follows. Section 2 formalizes the problem of learning the density for general probability measures and introduces our RKHS-based estimator. Section 3 presents theoretical guarantees, including consistency and finite-sample bounds. Section 4 discusses the low-rank approximation approach and its computational advantages. In Section 5, we apply our method to empirical tasks in domain adaptation, anomaly detection, causal inference, and high-dimensional conditional distribution estimation. Section 6 illustrates the applicablity of KDM in hypothesis testing and conditional distribution estimation for both simulated and real data. Section 7 concludes and points out future directions. The appendix contains additional results and all proofs.
2 General Density Estimation Problem
@@
We denote the density by $g_{\star}\coloneqq\frac{d{\mathbb{Q}}}{d{\mathbb{P}}}$.
The goal of this paper is to learn the true density $g_{\star}$ under minimal assumptions from samples of ${\mathbb{P}}$ and ${\mathbb{Q}}$. To this end, we assume that
@@
see, e.g., Steinwart and Scovel (2012, Section 2). Given (2), we can represent the hypothesis density as
@@
The additive splitting (4) is motivated by the fact that $g_{\star}$ itself might not be an element of ${\mathcal{H}}$ while $g_{\star}-p_{\star}$ is. To see the relevance of this approach, consider the example where ${\mathcal{Z}}={\mathbb{R}}$, $g_{\star}=1$ and $k$ is a Gaussian kernel. Indeed, this Gaussian RKHS does not contain the constant functions (in fact, it contains no polynomials, see Minh (2010)). In this example, one could thus choose the constant prior $p_{\star}=1$. In general, we neither assume that $p_{\star}$ is positive nor that it integrates to 1 with respect to ${{\mathbb{P}}}$. Hence, $p_{\star}=0$ is a possible choice. Notably, our specification of the hypothesis density (4) is novel to the literature, as it is more flexible and requires only the milder assumption (2), in contrast to the stronger assumptions typically employed in previous work. For a more detailed comparison, see Section 5.2.3 below.
@@
The error function ${\mathcal{E}}(h)$, and the expression in (8), contain the unobservable population density $g_{\star}$ and therefore are not readily available for estimations based on samples of ${\mathbb{P}}$ and ${\mathbb{Q}}$. We solve this problem by using the identity $J_{\mathbb{P}}^{\ast}g_{\star}=J_{\mathbb{Q}}^{\ast}1$ justified by the following lemma.
@@
To assess how close $p_{\star}+J_{\mathbb{P}}h_{\lambda}$ is to the true density $g_{\star}$, we consider the orthogonal decomposition of the squared population error
@@ 4 Low-rank Approximation
The following classic result asserts that Algorithm 1 works, see, e.g., Horn and Zhang (2005), Harbrecht et al. (2012) or Filipovic et al. (2025). A novelty of Algorithm 1 lies in the computation of the biorthogonal matrix $\bm{R}$ at no additional complexity along with $\bm{L}$. For completeness, we provide a self-contained proof in Appendix A.
@@ 5.1 Hypothesis Testing
As a first application, we obtain statistical tests of the hypothesis that the true equals the prior density,
@@
For $p_{\star}=1$, the hypothesis (38) is equivalent to ${\mathbb{Q}}={\mathbb{P}}$, which is the same hypothesis as studied in Pfister et al. (2017). Then the sample variable $u_{\lambda}$ in (82) in the proof of Theorem 5.1 is the scaled difference $S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{1}$ between the sample kernel mean embeddings of ${\mathbb{Q}}$ and ${\mathbb{P}}$. As a corollary of Theorem 5.1 we thus recover the kernel two-sample test of Gretton et al. (2012) as a special case, connecting their maximum mean discrepancy approach (Gretton et al., 2012, Theorem 5) with the Hilbert–Schmidt independence criterion (Pfister et al., 2017, Proposition 1).
@@ 5.2.1 Derivation of the Estimator
For any hypothesis density (4) with $h\in{\mathcal{H}}$, we thus obtain the corresponding model of the true conditional expectation ${\mathbb{E}}[f(Y)\mid X=x]=\int_{\mathcal{Y}}f(y){\mathbb{P}}_{Y\mid X=x}(dy)$ given by
@@
for any $f\in L^{2}_{{\mathbb{P}}_{Y}}$. Here we estimate the ${\mathbb{P}}_{Y}$-integral in (49) by the empirical counterpart given by (50), for an auxiliary i.i.d. sample $\bar{y}_{1},\dots,\bar{y}_{\bar{n}}$ of ${\mathbb{P}}_{Y}$. The last approximation in (51) is practically motivated such that we integrate $f$ with respect to a bona-fide conditional distribution. [^13] The density $g_{\star}$ satisfies the implicit structural properties $g_{\star}(x,y)\geq 0$ ${\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$-a.s., $\int_{\mathcal{Y}}g_{\star}(x,y){\mathbb{P}}_{Y}(dy)=1$ ${\mathbb{P}}_{X}$-a.s., and $\int_{\mathcal{X}}g_{\star}(x,y){\mathbb{P}}_{X}(dx)=1$ ${\mathbb{P}}_{Y}$-a.s.
The estimator $\widetilde{h}_{\lambda}$ in (49) requires i.i.d. samples $z_{{\mathbb{P}},1},\dots,z_{{\mathbb{P}},n}$ of ${\mathbb{P}}={\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$ and $z_{{\mathbb{Q}},1},\dots,z_{{\mathbb{Q}},n}$ of ${\mathbb{Q}}={\mathbb{P}}_{(X,Y)}$. If we could only sample from the joint distribution, we would consider an i.i.d. sample $(x_{1},y_{1}),\dots,(x_{3n},y_{3n})$ of ${\mathbb{P}}_{(X,Y)}$ of size $3n$ and set
@@ 5.2.3 Comparison with Related Literature
In view of the decomposition (48) it is natural to compare conditional expectations (46) calculated from a hypothesis density (4) to the conditional mean embedding introduced in Song et al. (2009). To facilitate this comparison, we assume in this subsection that ${\mathcal{H}}={\mathcal{H}}_{\mathcal{X}}\otimes{\mathcal{H}}_{\mathcal{Y}}$ is the tensor product of two separable RKHS ${\mathcal{H}}_{\mathcal{X}}$ and ${\mathcal{H}}_{\mathcal{Y}}$ with measurable kernels $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$ on ${\mathcal{X}}$ and ${\mathcal{Y}}$, respectively. In line with (3), we also assume that $\sup_{x\in{\mathcal{X}}}k_{\mathcal{X}}(x,x)<\infty$ and $\sup_{y\in{\mathcal{Y}}}k_{\mathcal{Y}}(y,y)<\infty$, so that the canonical embeddings $J_{{\mathbb{P}}_{X}}\colon{\mathcal{H}}_{\mathcal{X}}\to L^{2}_{{\mathbb{P}}_{X}}$, $J_{{\mathbb{P}}_{Y}}\colon{\mathcal{H}}_{\mathcal{Y}}\to L^{2}_{{\mathbb{P}}_{Y}}$ are Hilbert–Schmidt operators. We then have $k\big{(}(x_{1},y_{1}),(x_{2},y_{2})\big{)}=k_{\mathcal{X}}(x_{1},x_{2})k_{\mathcal{Y}}(y_{1},y_{2})$, and in view of (47) we can identify the operators $J_{{\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}}=J_{{\mathbb{P}}_{X}}\otimes J_{{\mathbb{P}}_{Y}}$.
@@
Our approach, via the density $g_{\star}$, overcomes the above limitations of the traditional conditional mean embedding approach. First, our construction of $\mu_{Y\mid X=x}$ is more elementary and feasible under verifiable technical assumptions, such as (1) and (2). Second, our hypothesis model (4) of $g_{\star}$ yields the candidate conditional distribution ${\mathbb{P}}_{Y\mid X=x}$ directly using identity (45). Third, we thus compute conditional expectations (46) directly for all test functions $f\in L^{2}_{{\mathbb{P}}_{Y}}$, and not only for $f\in{\mathcal{H}}_{\mathcal{Y}}$.
@@
Also the setup in Suzuki et al. (2009); Sugiyama (2013) is restricted to continuous variables with a Lebesgue density. In fact, they assume the assumptions of Lemma 5.3 hold for the Lebesgue measures $\mu_{\mathcal{X}}(dx)=dx$ and $\mu_{\mathcal{Y}}(dy)=dy$. In this case, the “density $w(x,y)$” in Suzuki et al. (2009) is equal to our density $g_{\star}(x,y)$. Their estimator of the density is based on a sampling scheme, similar to Filipovic et al. (2025), which is different from ours. As a consequence, they do not provide finite sample guarantees nor a central limit theorem.
There is a vast literature on (semi-)parametric distributional regression approaches, which includes structured additive models, see Rügamer et al. (2023) for a recent overview. In contrast, our approach is nonparametric functional analytic and learns the density in a reproducing kernel Hilbert space, which comes with asymptotic theory and finite sample guarantees. Isotonic distribution regression in Henzi et al. (2021) is to date specialized on univariate response variables and conditional distributions, conformal predictive distributions from Vovk et al. (2019) are based on regression problems. Shen and Meinshausen (2024) estimates conditional distributions only.